Please help me with this question. A rectangle platform is 11 feet longer than it is wide. If the area of the platform is 276 sq. ft. find the length and the width. Please show all work necessary to arrive to the solution.
Answer by stanbon(75887)

Please help me with this question.
A rectangle platform is 11 feet longer than it is wide.
---
Let the width be "x"
Then the length is "x+11"
--------------------------------
If the area of the platform is 276 sq. ft. find the length and the width. Please show all work necessary to arrive to the solution.
--------------------------------
Equation:
Area = length * width
276 = (x+11)x
Rearrange:
x^2 + 11x - 276 = 0
Factor:
(x-12)(x+23) = 0
---
Positive solution:
x = 12 ft (width)
x+11 = 23 ft (length)
==========================
Cheers,
Stan H.
==========================